# egoPPG: Heart Rate Estimation from Eye-Tracking Cameras in Egocentric Systems to Benefit Downstream Vision Tasks

### v1

@@ Abstract
Egocentric vision systems aim to understand the spatial surroundings and the wearer’s behavior inside it, including motions, activities, and interaction with objects. Since a person’s attention and situational responses are influenced by their physiological state, egocentric systems must also detect this state for better context awareness. In this paper, we propose egoPPG, a novel task for egocentric vision systems to extract a person’s heart rate (HR) as a key indicator of the wearer’s physiological state from the system’s built-in sensors (e.g., eye tracking videos). We then propose EgoPulseFormer, a method that solely takes eye-tracking video as input to estimate a person’s photoplethysmogram (PPG) from areas around the eyes to track HR values—without requiring additional or dedicated hardware. We demonstrate the downstream benefit of EgoPulseFormer on EgoExo4D, where we find that augmenting existing models with tracked HR values improves proficiency estimation by 14%. To train and validate EgoPulseFormer, we collected a dataset of 13+ hours of eye-tracking videos from Project Aria and contact-based blood volume pulse signals as well as an electrocardiogram (ECG) for ground-truth HR values. 25 participants performed diverse everyday activities such as office work, cooking, dancing, and exercising, which induced significant natural motion and HR variation (44–164 bpm). Our model robustly estimates HR (MAE=8.82 bpm) and captures patterns (r=0.81). Our results show how egocentric systems may unify environmental and physiological tracking to better understand user actions and internal states.
Figure 1: Implementing a novel vision task egoPPG, EgoPulseFormer tracks a person’s heart rate (HR) from eye-tracking videos captured by unmodified egocentric vision headsets. Our method estimates the person’s photoplethysmogram (PPG) of the blood volume pulse from areas around the eyes to extract HR values. For training and validation, we collected egoPPG-DB, a dataset of participants’ eye-tracking videos during everyday activities with synchronized ground-truth PPG signals (white contact sensor) and HR values (ECG chest strap).
@@ 1 Introduction
Egocentric vision systems, such as Mixed Reality (MR) glasses by Meta [56], Magic Leap [46], and others have emerged as powerful devices for capturing and analyzing a person’s behavior as well as their surrounding environment from a first-person perspective. The wider availability of promising wearable capture platforms (e.g., Project Aria glasses [23]) has sparked a large amount of research on egocentric vision tasks for environment understanding and navigation, including localization [71, 41, 74], and simultaneous localization and mapping (SLAM) [17, 38, 68]. Since egocentric systems can simultaneously capture parts of the wearer’s behavior, many efforts have investigated egocentric action recognition [52, 89, 94, 44, 86] and hand-object interaction [73, 32, 95, 22] to understand user behavior. Large-scale datasets have also been introduced in recent years to accelerate data-driven research in this domain, such as Ego4D [32], Nymeria [51], and EgoExo4D [33], offering rich, multimodal data to enable training and evaluation for these tasks.
In addition to spatial awareness, understanding the user’s behavior, their attention, and intent is equally important for egocentric systems [88, 3, 58]. For instance, anticipating the user’s next action is crucial for applications in navigation, personalized feedback, and autonomous assistance [20, 91, 84]. Objects of interest are commonly estimated from analyzed gaze patterns [27, 47, 37, 45] to support behavioral analysis and social understanding [26, 40, 32, 33].
@@
However, holistically modeling a person’s behavior and intent requires knowledge of their physiological state, which influences cognitive performance, attention, and situational responses [25, 15, 53, 81, 76]. Key components of physiological state include cardiovascular indicators such as heart rate (HR) and electrodermal activity [59, 79, 1, 9], which reflect emotions, stress, fatigue, and alertness [65, 11, 59, 79, 1]. Capturing these dynamics can thus benefit models of human behavior to enable a richer understanding of user actions for adaptive systems.
In this paper, we introduce egoPPG, a novel task for egocentric vision systems to accurately extract a person’s HR measurements from the system’s built-in sensors, specifically the eye-tracking cameras in unmodified headsets. We then propose EgoPulseFormer, a novel method that implements egoPPG on Project Aria glasses to demonstrate the benefit of egoPPG for existing downstream vision tasks on large egocentric datasets. Our learning-based method EgoPulseFormer is designed to recover the person’s photoplethysmogram (PPG) from the subtle fluctuations in skin intensity due to light absorption in pulsatile arteries beneath the skin surface following a blood volume pulse (BVP), in particular deriving it from regions around the wearer’s eye for robust tracking. Leveraging the infrared (IR) illuminant of the eye tracker, it work across light conditions [50]. In addition, we demonstrate the benefits of estimated HR values in egocentric recordings for a key vision task downstream: We augment an existing architecture with EgoPulseFormer and show its impact on EgoExo4D’s proficiency estimation benchmark, whose accuracy improves by 14.1% with EgoPulseFormer’s HR estimates.
The sensing configuration of our task can be considered a hybrid between typical contact-based BVP sensors (e.g., those in smartwatches [5, 31, 63, 21, 13, 61]) and rPPG methods that aim to extract HR from a person’s face using a camera [39, 82, 66]. While egocentric glasses are body-worn much like wrist watches, they couple more loosely to the body and are subject to considerable motion artifacts. Unlike contact sensors, eye trackers observe the wearer’s eye regions from a short distance and capture eye motions and blinks, leading to ambiguity and noise for capturing fluctuations in skin intensity. Unlike rPPG configurations, egocentric capture systems move with the wearer’s body and head, and eye trackers use controlled illumination.
@@
Therefore, we designed EgoPulseFormer to extract BVP from the region with the least motion—around the wearer’s eyes—by incorporating spatial attention within our model’s backbone. We validate EgoPulseFormer’s efficacy on a novel dataset that we collected to capture some of the activities included in large-scale egocentric datasets alongside physiological reference recordings. Our dataset egoPPG-DB contains 13 hours of recordings from 25 participants, who wore Project Aria glasses and performed five real-world tasks with varying motion and intensity, causing their HR values to reach levels between 44–164 bpm.
@@
    egoPPG as a novel task and EgoPulseFormer as an HR estimation method for egocentric systems that operates on eye-tracking videos. Our method robustly predicts continuous HR across a series of activities and interactions (MAE=8.82 bpm), with a 27% lower error than current state-of-the-art rPPG models [12, 92, 48, 90].
@@
    egoPPG-DB, a dataset of eye-tracking videos and synchronized BVP (contact-based) and ECG recordings (chest strap-based) to verify all physiological signals. We captured these across diverse everyday activities that were inspired by those included in existing large-scale egocentric datasets [32, 51, 33].
@@
    a validation of egoPPG’s downstream benefits for proficiency estimation. Augmenting EgoExo4D with continuous HR values and additionally feeding them into an existing architecture, we demonstrate the implications of our method on the proficiency estimation benchmark with an increase in accuracy by 14.1%.
@@ 2 Related work
Egocentric vision. In recent years, research in egocentric vision has surged, driven by advances in AR/VR glasses [4, 36, 46, 56, 57, 23], which provide new ways for understanding user interaction from a first-person perspective. Much of this work has focused on tasks such as action recognition [52, 89, 94, 44, 86] and anticipation [30, 16, 86, 60], full-body pose estimation [75, 87], responding to user needs [69, 67, 91], and social behavior analysis [26, 40, 32]. Additionally, tracking vital signs in AR/VR settings and for affective computing applications [65, 11, 59, 79, 1] has become an important tool for understanding users’ physiological states [54], further aiding in understanding users’ behavior, their attention, and intent [88, 3, 58].
Physiological measurements. Wearable sensors have had a tremendous impact on health monitoring in recent years, enabling continuous measurement of key physiological metrics, such as heart rate (HR), oxygen saturation, and activity levels [63, 21, 13, 62]. Heart rate (HR), in particular, is a key measure for assessing an individual’s health and performance [29, 24, 72, 42]. While wearable sensors, such as wrist-worn smartwatches, provide accurate HR measurements and also challenging scenarios (e.g., exercising), they are intrusive and can cause discomfort [43]. Recent research has, thus, extensively explored using cameras as an unobtrusive, non-contact alternative for measuring HR, generally called remote photoplethysmography (rPPG) [39, 82, 66]. rPPG measures HR based on subtle color changes in the skin caused by the BVP. Generally, rPPG methods can be broadly divided into traditional signal processing techniques [39, 82, 66, 8, 18, 19, 83] and deep learning-based approaches [12, 92, 48, 10, 90]. So far, rPPG has been mostly applied to facial videos with the camera and user being stationary, such as while sitting in front of a laptop, as it requires a continuous video feed of the same skin region. This limitation is shown in current rPPG datasets, which primarily capture individuals in seated positions with either a stationary camera directed at their face [78, 34, 7, 64, 70] or requiring users to hold a smartphone steadily in front of their face [80]. As a result, rPPG is not feasible to be deployed in more dynamic settings, such as during exercise.
Eye tracking cameras. Eye tracking in egocentric vision systems is mostly done using inward-facing cameras direct at the eyes [2]. Even during motion, eye tracking in VR devices demonstrated accurate performance showcasing that the cameras remain almost stationary relative to the user’s eyes [14]. Furthermore, IR illumination makes them robust to lighting variations and low-light conditions [50]. To the best of our knowledge, videos from eye tracking cameras have not yet been explored for HR estimation using the BVP despite their promise to enable unobtrusive HR measurements during everyday life.
@@ 3 Overview
Our aim is to enable egocentric vision systems i) to track a person’s physiological state via continuously estimated HR and ii) to integrate these HR estimates into downstream tasks that benefit from knowledge of the user’s state. Sec. 4 first describes our dataset of synchronized eye-tracking videos and ground-truth HR measurements during a series of everyday activities. Sec. 5 then outlines our method EgoPulseFormer that continuously estimates a person’s BVP from eye-tracking videos and derives HR values from it. Fig. 3 illustrates our approach. To demonstrate EgoPulseFormer’s benefit for downstream applications, we leverage HR estimation for the user proficiency benchmark of the EgoExo4D dataset, described in Sec. 6. Finally, Sec. 7 provides all results from our evaluations and Sec. 8 discusses our findings.
@@
The egoPPG-DB dataset was developed to support HR estimation from eye-tracking videos under real-world conditions, with a protocol designed to elicit significant motion and fluctuations in HR. By including diverse everyday activities, we provide a challenging benchmark for egocentric HR estimation models.
@@ 4.1 Recruiting and recording
We recruited $N=25$ participants (12 female, 13 male, ages 19–32, $\mu=25.1$ and $\sigma=3.3$) on a voluntary basis, resulting in over 13  hours of video recordings. Based on the Fitzpatrick scale [28], 9 participants had skin type II, 10 had skin type III, 2 had skin type IV, and 4 had skin type V. All participants signed a consent form before the data collection, agreeing with using and sharing their data for academic and non-commercial purposes. Participants were instructed to avoid wearing makeup prior to the recording. The data collection was approved by the anonymized Ethics Commission (no. #anonymized). In terms of dataset size by duration, egoPPG-DB is third amongst the longest rPPG datasets [78, 34, 7, 64, 70, 80] as listed in Tab. 6.
@@ 4.3 Capture protocol
The average duration of the recording of participant was 32 minutes. The capture protocol comprised 5 activities (Tab. 1): watching a video, office work, kitchen work, dancing, and exercising on an indoor bike (Fig. 1). We included these activities for three purposes: (1) Incorporate everyday activities including the corresponding HR changes and motion artifacts; (2) cover a wide range of HR values (low HR when watching a video vs. high HR when exercising), and (3) resemble activities that were captured in large-scale egocentric vision datasets, such as EgoExo4D [33] or Nymeria [51]. In Tab. 8, we give a detailed description of each activity, and in Tab. 3 we show the mean HR values for each activity. Exercising on the bike produced the highest HR values (113 bpm), whereas watching the video resulted in the lowest mean HR (71 bpm).
@@ 4.4 Dataset and signal quality verification
To evaluate that the contact PPG sensor, whose signal we later use as the target for model training, produces accurate HR values, we evaluate it against the gold-standard ECG. We assessed the performance by calculating the MAE and Pearson correlation between HR estimates from the ECG and PPG signals for each participant using a 30-second sliding window. For activity labeling, we manually annotated the start and end times of each task (see Tab. 1) for each participant using the Point of View (POV) RGB videos recorded by the Project Aria glasses. To ensure that the signal quality of the contact PPG is sufficient for model training, we excluded all tasks with an MAE over 3.0 bpm between the PPG and ECG, which can happen when the PPG sensor occasionally loses alignment with the angular artery due to movement. This applied to 20 of 150 tasks (13%, see Tab. 7). During the remaining tasks, our custom-built PPG nose sensor achieved very high accuracy, with an average MAE of 1.3 bpm and a mean correlation of 0.94 compared to the ECG signal, showing its suitability as ground truth.
@@
5 EgoPulseFormer: a first method for egoPPG
@@ 5.1 Problem definition
Our objective is to estimate BVP and HR from periodic changes in pixel intensity in eye-tracking video frames $\boldsymbol{F}\in\mathbb{R}^{w\times h}$. Physically, this means extracting a physiological signals from the information in the light reflected by the arteries and arterioles that carry blood beneath the skin. This light reflection can be modeled as a combination of diffuse and specular reflections. Wang et al. [83] model the reflected light intensity $C(t)$ as:
@@
Figure 3: Architecture of our model for continuous BVP estimation from eye-tracking videos and consecutive HR computation.
@@ 5.2 Deep learning model
Our architecture is built upon a 3D CNN backbone [92] with a temporal input length of $T=128$ frames (corresponding to 4.3 seconds) downsampled to $(h=48)\times(w=128)$ pixels, resulting in an input of dimensions $(T,C,w,h)$. The channel is $C=1$ in our case, as our input is from monochrome videos. Compared to the video of a person’s face, which is usually used for rPPG tasks, eye tracking videos offer additional challenges. While the bulbar conjunctiva (white of the eyes) contains many blood vessels from which the BVP could theoretically be estimated, eyes typically move strongly during everyday situations and are closed while blinking (see participant 2 in Fig. 4). Consequently, extracting the BVP from the eye regions would introduce substantial motion artifacts and reduce the signal-to-noise ratio (SNR). In contrast, when qualitatively analyzing eye tracking images, we see that the skin around the eyes exhibits considerably less motion than the eyes themselves and could thus provide a more stable source of BVP information. To address this, we introduce spatial attention modules [85] before each pooling (see Fig. 3) to allow our network to focus on high-SNR regions, such as the skin, and reduce the influence of low-SNR regions with frequent motion, like the eyes. Given an input feature map $\boldsymbol{F}\in\mathbb{R}^{T\times C\times w\times h}$, the spatial attention modules infers a spatial attention map $\boldsymbol{M_{s}}\in\mathbb{R}^{T\times 1\times w\times h}$ as:
@@
where $\sigma$ is the sigmoid function, $f^{7\times 7}$ a $7\times 7$ convolution operation and $\boldsymbol{F_{avg}}\in\mathbb{R}^{T\times 1\times w\times h}$ and $\boldsymbol{F_{max}}\in\mathbb{R}^{T\times 1\times w\times h}$ are the average-pooled and max-pooled feature maps respectively. The final output $\mathbf{F_{out}}$ of the attention process is then:
$\mathbf{F_{out}}=\mathbf{M_{s}}\otimes\mathbf{F},$ (4)
Furthermore, individual variations in the fit of the glasses result in different parts of the skin around the eyes being visible. For some individuals, the eye tracking cameras capture only the areas above the eyes, for others, only below, and in some cases, the glasses sit at an incline (see Fig. 4). To account for such variations, we apply three targeted data augmentations during training that reflect these specific differences in camera angles and coverage: (1) random rotation between -20 and +20 degrees to account for slight inclinations in the glasses’ positioning; (2) random horizontal cropping to help the network distinguish between high and low SNR regions across various skin areas and camera positions; and (3) horizontal and vertical flipping to further increase robustness to individual differences in skin region visibility. Finally, to help the network focus on the changes between the frames caused by the periodic BVP, we use the standardized consecutive frame differences of the eye tracking videos as input into our network. We standardize each frame by subtracting the mean pixel intensity and dividing it by the standard deviation of the pixel intensities values [12]. We use the standardized consecutive differences of the PPG signals from the nose as the labels for our model. Fig. 3 shows our architecture with the input preprocessing demonstrated on the left side, and an example learned spatial attention map on the right side. For one batch, the total number of FLOPS for our model is about 164 GigaFLOPS. The total number of parameters is about 770k, where the spatial attention modules added only 400 parameters.
Figure 4: Left: Head geometry determines the regions that the eye tracker captures. Right: Learned spatial attention maps show that eye regions are excluded and EgoPulseFormer instead extracts BVP from the surrounding skin regions, which moves less than the eyes.

### v2

@@ Abstract
Egocentric vision systems aim to understand the spatial surroundings and the wearer’s behavior inside it, including motions, activities, and interactions. We argue that egocentric systems must additionally detect physiological states to capture a person’s attention and situational responses, which are critical for context-aware behavior modeling. In this paper, we propose egoPPG, a novel vision task for egocentric systems to recover a person’s cardiac activity to aid downstream vision tasks. We introduce PulseFormer, a method to extract heart rate as a key indicator of physiological state from the eye tracking cameras on unmodified egocentric vision systems. PulseFormer continuously estimates the photoplethysmogram (PPG) from areas around the eyes and fuses motion cues from the headset’s inertial measurement unit to track HR values. We demonstrate egoPPG’s downstream benefit for a key task on EgoExo4D, an existing egocentric dataset for which we find PulseFormer’s estimates of HR to improve proficiency estimation by 14%. To train and validate PulseFormer, we collected a dataset of 13+ hours of eye tracking videos from Project Aria and contact-based PPG signals as well as an electrocardiogram (ECG) for ground-truth HR values. Similar to EgoExo4D, 25 participants performed diverse everyday activities such as office work, cooking, dancing, and exercising, which induced significant natural motion and HR variation (44–164 bpm). Our model robustly estimates HR (MAE=7.67 bpm) and captures patterns (r=0.85). Our results show how egocentric systems may unify environmental and physiological tracking to better understand users and that egoPPG as a complementary task provides meaningful augmentations for existing datasets and tasks. We release our code, dataset, and HR augmentations for EgoExo4D to inspire research on physiology-aware egocentric tasks.
Figure 1: We propose egoPPG as a novel computer vision task: tracking a person’s heart rate (HR) on unmodified egocentric vision headsets. Taking eye tracking videos as input, our method PulseFormer estimates the photoplethysmogram (PPG) from areas around the eyes to derive HR values. For training and validation, we introduce egoPPG-DB, a dataset of eye tracking videos while participants performed everyday activities with synchronized ground-truth PPG (via nose-based contact sensor) and HR values (via ECG chest strap).
@@ 1 Introduction
Egocentric vision systems, such as Mixed Reality (MR) glasses by Meta [56], Magic Leap [43], and others have emerged as powerful devices for capturing and analyzing a person’s behavior and their environment from a first-person perspective. The wider availability of promising wearable capture platforms has sparked a large amount of research on egocentric vision tasks for environment understanding and navigation [18], including localization [71, 39, 74], and simultaneous localization and mapping [15, 33, 68]. Since egocentric systems simultaneously capture parts of the wearer’s behavior in addition to their environment, prior work has investigated egocentric action recognition [51, 89, 96, 87] and hand-object interaction [73, 26, 97] to understand user behavior. Several large-scale datasets now accelerate data-driven research in this domain with multimodal data for training and evaluation (e.g., Ego4D [26], Nymeria [50], EgoExo4D [27]).
@@
Most recently, Meta’s Project Aria 2 introduced a contact-based heart rate (HR) sensor, with which egocentric systems can gauge the wearer’s cognitive performance, attention, and situational responses [20, 13, 52, 83, 76]. Numerous additional conditions manifest in a person’s HR, such as emotions, stress and fatigue [65, 9, 59, 80, 1]—capturing these dynamics can thus benefit models of human behavior to enable a richer understanding of user behavior.
In this paper, we introduce a method to make such HR estimates available to many existing egocentric systems and already recorded large datasets, such as EgoExo4D [27] or Nymeria [50]. Our method PulseFormer accurately recovers a person’s HR from the eye tracking videos in egocentric headsets. PulseFormer first estimates the person’s photoplethysmogram (PPG) from the subtle fluctuations in skin intensity due to pulsatile artery expansion beneath the surface following a blood volume pulse (BVP), in particular deriving it from regions around the wearer’s eye for robust tracking. Our spatial attention module ensures that PPG is estimated from robust regions around the eye, while our cross-attention fusion with the system’s inertial measurement unit (IMU) learns a motion-informed temporal attention to optimally weight the eye tracking images for more accurate PPG estimates in scenarios with heavy motion.
@@
We validate our method’s efficacy on a novel dataset that we collected to capture some of the activities included in large-scale egocentric datasets alongside physiological reference recordings. Our dataset egoPPG-DB contains 13 hours of recordings from 25 participants, who wore Project Aria glasses and performed six real-world tasks with varying motion and intensity, causing their HR values to reach levels between 44–164 bpm.
Downstream benefits for egocentric vision tasks
A key contribution of our paper is that we demonstrate that knowing a person’s continuous HR values benefits egocentric vision tasks downstream. We augment an existing architecture with PulseFormer’s HR estimates and show its impact on EgoExo4D’s proficiency estimation benchmark, which improves accuracy on this task by 14.1%.
Contributions
@@
    egoPPG as a novel task and PulseFormer as an HR estimation method for egocentric systems that operates on eye tracking videos. Our method robustly predicts continuous HR across a series of activities and interactions (MAE=7.67 bpm), with a 23.8% lower error than current state-of-the-art rPPG models [10, 92, 45, 93, 90].
@@
    egoPPG-DB, a dataset of eye tracking videos and synchronized BVP (contact-based) and ECG recordings (chest strap-based) to verify all physiological signals. We captured these across diverse everyday activities that were inspired by those included in existing large-scale egocentric datasets, such as EgoExo4D [26, 50, 27].
@@
    a validation of egoPPG’s downstream benefits for egocentric vision tasks. We demonstrate the implications of our method PulseFormer on the proficiency estimation benchmark of the EgoExo4D dataset, which increases the accuracy by 14.1% when augmenting EgoExo4D with our continuously predicted HR values.
@@ 2 Related work
Egocentric vision. In recent years, research in egocentric vision has surged, driven by advances in AR/VR glasses [4, 31, 43, 56, 57, 18], which provide new ways for understanding user interaction from a first-person perspective. Much of this work has focused on tasks such as action recognition [51, 89, 96, 41, 87] and anticipation [25, 14, 87], full-body pose estimation [75, 36, 37], responding to user needs [69, 67, 91], and social behavior analysis [21, 35, 26]. Additionally, tracking vital signs in AR/VR settings and for affective computing applications [65, 9, 59, 80, 1] has become an important tool for understanding users’ physiological states [53], their behavior, attention, and intent [88, 3, 58].
Physiological measurements. Wearable sensors have had a tremendous impact on health monitoring in recent years, enabling continuous measurement of key physiological metrics, such as heart rate (HR), oxygen saturation, and activity levels [62, 17, 11, 61, 55]. HR, in particular, is a key measure for assessing an individual’s health and performance [23, 19, 72, 40, 48]. In parallel to wearable sensors such as smartwatches, recent research has extensively explored using cameras as an unobtrusive, non-contact alternative for measuring HR, generally called remote photoplethysmography (rPPG) [34, 84, 66]. rPPG measures HR based on the BVP via subtle color changes of the skin. Generally, rPPG methods can be broadly divided into traditional signal processing techniques [34, 66, 7, 16, 85] and deep learning-based approaches [10, 92, 45, 8, 90, 78]. So far, rPPG has been mostly applied to facial videos with the camera and user being stationary, such as while sitting in front of a laptop, as it requires a continuous video feed of the same skin region. This limitation is shown in current rPPG datasets, which primarily capture individuals in seated positions with either a stationary camera directed at their face [79, 29, 6, 70] or requiring users to hold a smartphone steadily in front of their face [82]. As a result, rPPG is not feasible to be deployed in more dynamic settings.
Eye tracking cameras. Eye tracking in egocentric vision systems is mostly done using inward-facing cameras directed at the eyes [2]. Even during motion, eye tracking in VR devices demonstrated accurate performance showcasing that the cameras remain almost stationary relative to the user’s eyes [12]. Furthermore, IR illumination makes them robust to lighting variations and low-light conditions [49]. To the best of our knowledge, videos from eye tracking cameras have not yet been explored for HR estimation.
@@ 3 Overview
Our aim is to enable egocentric vision systems i) to model a person’s physiological state via continuously estimated HR and ii) to integrate these HR estimates into downstream tasks that benefit from knowledge of the user’s state. Sec. 4 first describes our dataset of synchronized eye tracking videos and ground-truth HR measurements. Sec. 5 introduces our method PulseFormer for recovering continuous HR from eye tracking videos. Sec. 6 outlines the downstream benefits of our novel task, using HR as input for modeling user proficiency on the EgoExo4D dataset.
@@
The egoPPG-DB dataset was developed to support HR estimation from eye tracking videos under real-world conditions and contains significant motion and HR fluctuations. By including diverse everyday activities, we provide a challenging benchmark for egocentric HR estimation models.
@@ 4.1 Recruiting and recording
We recruited $N=25$ participants (12 female, 13 male, ages 19–32, $\mu=25.1$ and $\sigma=3.3$) on a voluntary basis, resulting in over 13  hours of video recordings. Based on the Fitzpatrick scale [22], 9 participants had skin type II, 8 had skin type III, 3 had skin type IV, and 5 had skin type V. All participants signed a consent form before the data collection, agreeing with using and sharing their data for academic and non-commercial purposes. The data collection was approved by the ETH Zurich Ethics Commission (no. 2023-N-08). In terms of duration, egoPPG-DB is among the longest rPPG datasets as listed in Tab. 10 (Supplementary). Participants of egoPPG-DB are not included in EgoExo4D.
@@ 4.3 Capture protocol
The average recording lasted 32 minutes. The capture protocol comprised 5 activities (Tab. 1): watching a video, office work, kitchen work, dancing, and exercising on an indoor bike (Fig. 1). We included these activities for three purposes: (1) Incorporate everyday activities including the corresponding HR changes and motion artifacts; (2) cover a wide range of HR values (low HR when watching a video vs. high HR when exercising), and (3) resemble activities that were captured in large-scale egocentric vision datasets, such as EgoExo4D [27]. In Tab. 11 (Supplementary), we give a detailed description of all activities. Tab. 3 shows mean HR values for each activity. Exercising on the bike produced the highest mean HR values (113 bpm), whereas watching the video resulted in the lowest (71 bpm).
@@ 4.4 Dataset and signal quality verification
To ensure the contact PPG sensor, whose signal we later use as the target for model training, produces accurate HR values, we evaluate it against the gold-standard ECG. We calculated the MAE and Pearson correlation between HR estimates from the ECG and PPG signals for each participant using a 30-second sliding window. For activity labeling, we manually annotated the start and end times of each task (see Tab. 1) for each participant using the Point of View (POV) RGB videos recorded by the Project Aria glasses. To ensure that the signal quality of the contact PPG is sufficient for model training, we excluded all tasks with an MAE over 3.0 bpm between the PPG and ECG (e.g. when the PPG sensor moved). This applied to 20 of 150 tasks (13%, see Tab. 6 in Supplementary). During the remaining tasks, our custom-built PPG nose sensor achieved very high accuracy, with an MAE of 1.3 bpm and a correlation of 0.94 compared to the ECG signal, showing its suitability as ground truth.
@@
5 PulseFormer method
@@ 5.1 Problem definition
Our objective is to estimate BVP and HR from periodic changes in pixel intensity in eye tracking video frames $\boldsymbol{F}\in\mathbb{R}^{w\times h}$. Physically, this means extracting physiological signals from the information in the light reflected by the arteries and arterioles that carry blood beneath the skin. This light reflection can be modeled as a combination of diffuse and specular reflections. Wang et al[85] model the reflected light intensity $C(t)$ as:
@@
Figure 3: Architecture of our model for continuous BVP estimation from eye tracking videos and consecutive HR computation.
@@ 5.2 Deep learning model
Our architecture is built upon a 3D CNN backbone (PhysNet) [92] with a temporal input length of $T=128$ frames (corresponding to 4.3 seconds) downsampled to $(h=48)\times(w=128)$ pixels, resulting in an input of dimensions $(T,C,w,h)$. The channel is $C=1$ in our case, as our input is from monochrome videos. The input in our network is the consecutive standardized frame differences (per participant frame-wise differences divided by the STD of the frames) of the eye tracking videos to help the network focus on the changes between frames [10]. As labels, we use the standardized consecutive differences of the PPG signals. Since eye tracking videos offer additional challenges compared to facial videos, usually used for rPPG tasks, we have designed our model to address these challenges (see Fig. 3).
Motion-informed temporal attention (MITA). Egocentric glasses are body-worn and subject to considerable motion artifacts when the user moves. Therefore, we propose to leverage the IMU within the glasses to obtain a motion-informed temporal attention. We employ a cross-attention module to integrate the IMU data with the video input, allowing our model to weigh each frame differently along the temporal dimension based on the motion intensity encoded by the IMU. This allows the model, e.g., to give less emphasis to frames heavily affected by motion. Given the input feature map $\boldsymbol{F_{in}}\in\mathbb{R}^{T\times 1\times w\times h}$, we use ResNet18 [28] and a linear layer to obtain the image embeddings $\boldsymbol{F_{e}}\in\mathbb{R}^{T\times D}$, where $D=128$ is the embedding dimension. Given IMU measurements $\mathbf{I_{in}}\in\mathbb{R}^{T\times 1}$, we use two 1D convolutional layers to obtain the IMU embeddings $\mathbf{I_{e}}\in\mathbb{R}^{T\times D}$. We then calculate the cross-attention $\mathbf{A}\in\mathbb{R}^{T\times D}$ as:
$\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{D}}\right)\mathbf{V}$ (3)
where $\mathbf{I_{e}}$ serve as queries $Q$ and $\mathbf{F_{e}}$ as keys $K$ and values $V$. Using a linear layer, we obtain the motion-informed temporal attention $\mathbf{T}\in\mathbb{R}^{T\times 1\times 1\times 1}$, which we multiply with $\boldsymbol{F_{in}}$.
Spatial attention (SA). While the bulbar conjunctiva (white of the eyes) contains many blood vessels from which the BVP could theoretically be estimated, eyes typically move strongly during everyday situations and are closed while blinking (see participant 2 in Fig. 4). Consequently, extracting the BVP from the eye regions would introduce substantial motion artifacts and reduce the signal-to-noise ratio (SNR). In contrast, when qualitatively analyzing eye tracking images, we see that the skin around the eyes exhibits considerably less motion than the eyes themselves and could thus provide a more stable source of BVP information. To address this, we introduce spatial attention modules [86, 32, 64] before each pooling (see Fig. 3) to allow our network to focus on high-SNR regions, such as the skin, and reduce the influence of low-SNR regions with frequent motion, like the eyes. Given some feature map $\boldsymbol{F}\in\mathbb{R}^{T\times C\times w\times h}$, the spatial attention modules infer a spatial attention map $\boldsymbol{M_{s}}\in\mathbb{R}^{T\times 1\times w\times h}$ as:
@@
where $\sigma$ is the sigmoid function, $f^{7\times 7}$ a $7\times 7$ convolution operation and $\boldsymbol{F_{avg}}\in\mathbb{R}^{T\times 1\times w\times h}$ and $\boldsymbol{F_{max}}\in\mathbb{R}^{T\times 1\times w\times h}$ are the average-pooled and max-pooled feature maps respectively. The final output $\boldsymbol{F_{out}}\in\mathbb{R}^{T\times C\times w\times h}$ of each attention process is then the product of $\boldsymbol{M_{s}}$ and $\boldsymbol{F}$.
Data augmentation. Furthermore, individual variations in the fit of the glasses result in different parts of the skin around the eyes being visible. For some individuals, the eye tracking cameras capture only the areas above the eyes, for others, only below, and in some cases, the glasses sit at an incline (see Fig. 4). To account for such variations, we apply three targeted data augmentations during training that reflect these specific differences in camera angles and coverage: (1) random rotation between -20 and +20 degrees to account for slight inclinations; (2) random horizontal cropping to help the network distinguish between high and low SNR regions across various skin areas and camera positions; and (3) horizontal and vertical flipping to further increase robustness to differences in skin region visibility. Our model requires approximately 399 GFLOPS per batch and has about 12M parameters. The frame rate is 2.9k fps on an RTX 4090 and 180 fps on an AMD EPYC CPU.
Figure 4: Left: Head geometry determines the regions that the eye tracker captures. Right: Learned spatial attention maps show that eye regions are excluded and PulseFormer instead extracts BVP from the surrounding skin regions, which moves less than the eyes.